Lean Accounting

Com-Tel Inc. manufactures and assembles two models of smartphones-the Tiger Model and the Lion Model. The process consists of a lean cell for each product. The data that follow concern only the Lion Model lean cell.

For the year, Com-Tel Inc. budgeted these costs for the Lion Model production cell:

Conversion Cost Categories Budget
Labor $69,700
Supplies 26,400
Utilities 9,500
   Total $105,600

Com-Tel plans 1,600 hours of production for the Lion Model cell for the year. The materials cost is $56 per unit. Each assembly requires 24 minutes of cell assembly time. There was no May 1 inventory for either Raw and In Process Inventory or Finished Goods Inventory.

The following summary events took place in the Lion Model cell during May:

  1. Electronic parts were purchased to produce 5,450 Lion Model assemblies in May.
  2. Conversion costs were applied for 5,200 units of production in May.
  3. 5,100 units were completed and transferred to finished goods in May.
  4. 4,950 units were shipped to customers at a price of $264 per unit.

If required, round your answers to the nearest cent.

Required:

1. Determine the budgeted cell conversion cost per hour.
$ per hour

2. Determine the budgeted cell conversion cost per unit.
$ per unit

3. Journalize the summary transactions (a) through (d).

a.
b.
c.
d. Sale
d. Cost

4. Determine the ending balance in Raw and In Process Inventory and Finished Goods Inventory.

Raw and In Process Inventory $
Finished Goods Inventory $

5. Lean accounting is different from traditional accounting because it is more   and uses   control. As a result, the number of transactions are  . In many lean operations, purchased materials are charged to a  . Direct labor is frequently  . Often, nonfinancial performance measures, such as  , are used to monitor performance.
Answer #1
1
Budgeted cell conversion cost per hour 66 per hour =105600/1600
2
Budgeted cell conversion cost per unit 26.40 per unit =66*(24/60)
3
Raw and in process inventory 305200 =5450*56
     Accounts Payable 305200
Raw and in process inventory 137280 =5200*26.40
       Conversion costs 137280
Finished goods inventory 420240 =5100*(56+26.40)
       Raw and in process inventory 420240
Accounts Receivable 1306800 =4950*264
       Sales 1306800
Cost of goods sold 407880 =4950*(56+26.40)
        Finished goods inventory 407880
4
Raw and In Process Inventory 22240 =305200+137280-420240
Finished Goods Inventory 12360 =420240-407880
5
Lean accounting is different from traditional accounting because it is more simplified and uses minimal control. As a result, the number of transactions are reduced . In many lean operations, purchased materials are charged to a “raw and in process inventory” account . Direct labor is frequently included as a conversion cost of the cell . Often, nonfinancial performance measures, such as lead time or quality measures, are used to monitor performance.